# Phantom Edges in the Problem Hamiltonian: A Method for Increasing Performance and Graph Visibility for QAOA

### v1

@@ IV Derivation of Max-Cut Expectation Value with $G^{\prime}$ Phase Operator
$\langle C_{uv}\rangle$ $=\frac{1}{2}+\frac{\sin(4\beta)}{4}\sin\gamma\left(\cos^{d}\gamma\cos^{d^{\prime}}(\alpha\gamma)+\cos^{e}\gamma\cos^{e^{\prime}}(\alpha\gamma)\right).$ (8)
@@ IV-A Example
In this section, we analytically demonstrate an improvement in the approximation ratio for a cycle graph, illustrating the utility of placing new edges in $G^{\prime}$ to form triangles with the original edges in $G$.
Figure 1: Examples of 8-vertex cycle graphs $G$ shown in red and additional edges from the complement graph $\overline{G}$ shown in green. The left image shows a $G^{\prime}$ that is still regular and triangle-free, while the right image shows a $G^{\prime}$ that is still regular and now contains triangles.
Let $G$ be an arbitrary-sized cycle graph. First, let’s consider optimizing the Max-Cut expectation value for $G$ using the standard QAOA ansatz. For a $D$-regular triangle-free graph, the expectation value is given by
@@
Now, let’s consider our modified ansatz with a phase operator reflecting a new graph $G^{\prime}$. If we add the $\alpha$-weighted edges to $G$ such that $G^{\prime}$ is triangle-free (left graph in Figure 1), then the QAOA cut expectation is symmetric across each edge in $G$. For the general case, where the cycle graph has any number of vertices and the $\alpha$-weighted edges connect all vertices that are three hops away, we arrive at the following formula. In this case, we apply Eq. 8 with parameters $d=e=1$ and $d^{\prime}=e^{\prime}=2$ for all edges. By summing the expectation $\langle C_{uv}\rangle$ over all edges in $G$, we obtain
@@
Next, consider a $G^{\prime}$ containing triangles (right graph in Figure 1). Again, due to symmetry, each edge has the same parameters $d,e$, and $f$. Substituting $d=e=1$, $d^{\prime}=e^{\prime}=2$, $f^{\prime}=1$, and $f=f^{\prime\prime}=0$ into Eq.  1 and summing over all edges for the general cycle graph gives:
@@
It is easy to see that the function is periodic with a maximum cut value of $0.7925m$. For even-lengthed cycles, this results in an improved approximation ratio of $\frac{0.7925m}{m}=0.7925$ for our modified ansatz, compared to the optimal approximation ratio of $\frac{0.75m}{m}=0.75$ for the standard QAOA ansatz. Note that greater Max-Cut values can be found if we optimize $\gamma,\beta$ along with $\alpha$.
@@ V-A Experimental Setup
For the following results, we compare two methods: the “Full Method” and the “Triangle Method.” As discussed in Section III-A, the Full Method is a straightforward approach that involves constructing a graph $G^{\prime}$ by adding an $\alpha$-weighted edge for every missing edge in the original graph $G$, making $G^{\prime}$ fully connected. In contrast, the Triangle Method only adds an $\alpha$-weighted edge between nodes in $G$ if they are two hops away, thereby creating triangles.
We applied both methods at depths $p=1$ and $p=2$ for nodes ranging from $6$ to $16$ and for all possible degrees [^1] It is an elementary result of graph theory that it is impossible to have an $k$-regular graph with $n$ vertices if both $n$ and $k$ are odd. for each node. The regular graphs were generated randomly using the Graphs.SimpleGraphs.random_regular_graph function in the Julia library Graphs.jl [7]. For each node and degree combination, we averaged the results over 10 non-isomorphic graphs whenever possible.
Figure 2: The average approximation ratios on 16-node graphs with degrees 4, 8, and 12 across various QAOA ansatzes: the standard ansatz at p=1 and p=2 (denoted ’p=1’ and ’p=2’) and our modified ansatz (denoted ’p=1 modified’ and ’p=2 modified’). For this plot, we constructed our modified phase operator using the Triangle Method.
@@ V-B Results
Fig.  2 compares the average approximation ratios between the standard ansatz and our modified ansatz at depths $p=1$ and $p=2$. We observe that for various levels of sparseness, the modified ansatz offers a significant improvement in the approximation ratio over the standard ansatz. Additionally, we note that for higher degrees (degree 12, in this instance), the $p=1$ modified ansatz performs as well as the standard $p=2$ QAOA. This similar trend has been observed across various numbers of nodes. For lower degrees (i.e., when the graph is more sparse), the $p=1$ modified ansatz yields approximation ratios that are approximately halfway between those of QAOA at $p=1$ and $p=2$.
As shown in the bar plot, the optimized approximation ratios generally cluster closely together, ranging from approximately 0.75 to 0.95. To facilitate easier interpretation, we plot the differences in approximation ratios. In Fig. 3 we show the average improvement in approximations ratio for various ansatzes for both the Full Method and Triangle Method at $p=1$ and $p=2$. We plot the average improvement as a function of graph degree $d$, with each line corresponding to a particular number of nodes $n$. As shown for both $p=1$ and $p=2$, the Triangle Method consistently performs as well as or better than the Full Method across all graph instances. For graphs with smaller degree, the Full Method yields worse approximation ratios compared to the Triangle Method, while for higher degrees, both methods yield similar approximation ratios.
When comparing the improvements for $p=1$ and $p=2$, we observe that the average improvement for $p=1$ is approximately $0.04$, while for $p=2$, it is around $0.02$. However, for $p=2$, there is a slight upward trend in improvement as the degree increases, a pattern consistent across all values of $n$. It is also important to note that although the average improvement decreases from $p=1$ to $p=2$, the baseline approximation ratios at $p=2$ are closer to $1$ (as seen by comparing the approximation ratios for the standard ansatz in  2), and therefore there will be less room for improvement at $p=2$.
We also plot the value of $\alpha$ that yields the maximum improvement in the approximation ratio, denoted as $\alpha_{\text{max}}$ in Fig.  4. The comparison between the Triangle Method and the Full Method follows a similar trend as for the approximation ratios: for higher degrees, $\alpha_{\text{max}}$ is the same for both methods. Additionally, we observe an upward trend in $\alpha_{\text{max}}$ as the degree increases, consistent across both methods and for both values of $p$.
@@ VI Discussion
To illustrate this effect on periodicity, in Fig. 5 we show a comparison of cost landscapes for the $12$-node cycle graph between the standard ansatz and our modifeid ansatz. We see that for the modified cost landscape that although the value of $\alpha=0.7$ is relatively high, there still exists peaks that yield an improved approximation ratio over the original ansatz as long as the parameter search spaces is expanded from $\gamma\in[-\pi,\pi]$, $\beta\in[-\pi/4,\pi/4]$ (green border) to $\gamma\in[-2\pi,2\pi]$, $\beta\in[-\pi/2,\pi/2]$ (blue border). On the other hand, when the search is restricted to the original parameter space, the peak energy values for the modified ansatz are lower than those for the standard ansatz. For practical purposes, we restrict our cost landscape to the region $\gamma\in[-\pi,\pi]$ and $\beta\in[-\pi/4,\pi/4]$, as optimizing over the entire possible region would be intractable. Also, the global optimum will on average occur near zero for non-periodic functions [18, 4, 5].
@@
One potential concern with our proposed QAOA ansatz is the number of two-qubit gates introduced into the circuit, as two-qubit gates are the primary source of error on modern NISQ devices. The number of two-qubit gates added is equal to the number of new edges in the intersection of $G^{\prime}$ and the complement of $G$. Therefore, the number of additional two-qubit gates depends on the original graph and the method used (e.g., Full Method or Triangle Method), providing some flexibility in our approach. Users can adjust the ansatz based on their specific needs. Additionally, if we were to increase $p$ by 1 instead, the number of additional two-qubit gates would equal the number of edges in $G$.
@@
Figure 5: Energy landscapes for (a) the standard QAOA ansatz and (b) the modified QAOA ansatz with a phase operator reflecting $G^{\prime}$, constructed by the Triangle Method with $\alpha=0.7$. The region within the green border is the region we explore in our numerical experiments. We see that greater peak energies can be found outside of the original search space.

### v2

@@ IV Derivation of Max-Cut Expectation Value for Phantom-QAOA
$\langle C_{uv}\rangle$ (10)
$=\frac{1}{2}+\frac{\sin(4\beta)}{4}\sin\gamma\left(\cos^{d}\gamma\cos^{d^{\prime}}(\alpha\gamma)+\cos^{e}\gamma\cos^{e^{\prime}}(\alpha\gamma)\right).$
@@ IV-A Example
In this section, we analytically show an improvement in the approximation ratio for a cycle graph. This example demonstrates the utility of adding new edges in $G^{\prime}$ such that additional triangles are formed with the existing edges in $G$.
Figure 2: Examples of 8-vertex cycle graphs $G$ shown in black and additional $\alpha$-weighted edges shown in red. The left image shows a $G^{\prime}$ that is regular and triangle-free, while the right image shows a $G^{\prime}$ that is also regular but contains triangles.
Let $G$ be an arbitrary-sized cycle graph. Let’s consider optimizing the Max-Cut expectation value for $G$ using the standard QAOA ansatz. For a $D$-regular triangle-free graph, the expectation value is given by
@@
Now, let’s consider the Phantom-QAOA ansatz with a phase operator reflecting a new graph $G^{\prime}$. If we add the $\alpha$-weighted edges to $G$ such that $G^{\prime}$ is triangle-free (left graph in Fig. 2), then the QAOA cut expectation is symmetric across each edge in $G$. For the general case, where the cycle graph has any number of vertices and the $\alpha$-weighted edges connect all vertices that are three hops away, we apply Eq. 10 with parameters $d=e=1$ and $d^{\prime}=e^{\prime}=2$ for all edges. By summing the expectation $\langle C_{uv}\rangle$ over all edges in $G$, we obtain
@@
Next, consider a $G^{\prime}$ containing triangles (right graph in Fig. 2). Again, due to symmetry, each edge has the same parameters $d,e$, and $f$. Substituting $d=e=1$, $d^{\prime}=e^{\prime}=2$, $f^{\prime}=1$, and $f=f^{\prime\prime}=0$ into Eq.  2 and summing over all edges for the general cycle graph gives
@@
It is easy to see that the function is periodic with a maximum cut value of $0.7925m$. For even-lengthed cycles, this results in an improved approximation ratio of $\frac{0.7925m}{m}=0.7925$ for Phantom-QAOA, compared to the optimal approximation ratio of $\frac{0.75m}{m}=0.75$ for the standard QAOA ansatz. Note that greater Max-Cut values can be found if we optimize $\gamma,\beta$ along with $\alpha$.
@@ V-A Experimental Setup
For the following results, we compare the two methods mentioned previously: the Full and Triangle methods. We applied both methods at depths $p=1$ and $p=2$ for nodes ranging from $6$ to $16$ and for all possible degrees [^1] It is an elementary result of graph theory that it is impossible to have an $k$-regular graph with $n$ vertices if both $n$ and $k$ are odd. for each node. The regular graphs were generated randomly using the Graphs.SimpleGraphs.random_regular_graph function in the Julia library Graphs.jl [7]. For each node and degree combination, we averaged the results over 10 non-isomorphic graphs whenever possible.
Figure 3: The average approximation ratios on 16-node graphs with degrees 4, 8, and 12 across various QAOA ansatzes: the standard ansatz at p=1 and p=2 (denoted ’p=1’ and ’p=2’) and the Phantom-QAOA ansatz (denoted ’p=1 Phantom’ and ’p=2 Phantom’). Here we used the Triangle method for Phantom-QAOA.
@@ V-B Results
Fig.  3 compares the average approximation ratios between the standard ansatz and our modified ansatz at depths $p=1$ and $p=2$. We observe that for various levels of sparseness, Phantom-QAOA offers a significant improvement in the approximation ratio over the standard ansatz. Additionally, we note that for higher degrees (degree 12, in this instance), the $p=1$ Phantom-QAOA ansatz performs as well as the standard $p=2$ QAOA. This similar trend was observed across various numbers of nodes. For lower degrees (i.e., when the graph is more sparse), the $p=1$ Phantom-QAOA ansatz yields approximation ratios that are approximately halfway between those of QAOA at $p=1$ and $p=2$.
As shown in the bar plot, the optimized approximation ratios generally cluster closely together, ranging from approximately 0.75 to 0.95. To facilitate easier interpretation, we plot the differences in approximation ratios. In Fig. 4 we show the average improvement in approximations ratio for various ansatzes for both the Full and Triangle methods at $p=1$ and $p=2$. We plot the average improvement as a function of graph degree $d$, with each line corresponding to a particular number of nodes $n$. As shown for both $p=1$ and $p=2$, the Triangle method consistently performs as well as or better than the Full method across all graph instances. For graphs with smaller degree, the Full method yields worse approximation ratios compared to the Triangle method, while for higher degrees, both methods yield similar approximation ratios.
When comparing the improvements for $p=1$ and $p=2$, we observe that the average improvement for $p=1$ is approximately $0.04$, while for $p=2$, it is around $0.02$. However, for $p=2$, there is a slight upward trend in improvement as the degree increases, a pattern consistent across all values of $n$. It is also important to note that although the average improvement decreases from $p=1$ to $p=2$, the baseline approximation ratios at $p=2$ are closer to $1$ (as seen by comparing the approximation ratios for the standard ansatz in  3), and therefore there will be less room for improvement at $p=2$.
We also plot the value of $\alpha$ that yields the maximum improvement in the approximation ratio, denoted as $\alpha^{*}$ in Fig.  5. The comparison between the Triangle and the Full methods follows a similar trend as for the approximation ratios: for higher degrees, $\alpha^{*}$ is the same for both methods. Additionally, we observe an upward trend in $\alpha^{*}$ as the degree increases, consistent across both methods and for both values of $p$.
@@ VI Discussion
To illustrate this effect on periodicity, in Fig. 6 we show a comparison of cost landscapes for the $12$-node cycle graph between the standard ansatz and Phantom-QAOA. We see that for the modified cost landscape that although the value of $\alpha=0.7$ is relatively high, there still exists peaks that yield an improved approximation ratio over the original ansatz as long as the parameter search spaces is expanded from $\gamma\in[-\pi,\pi]$, $\beta\in[-\pi/4,\pi/4]$ (green border) to $\gamma\in[-2\pi,2\pi]$, $\beta\in[-\pi/2,\pi/2]$ (blue border). On the other hand, when the search is restricted to the original parameter space, the peak energy values for the Phantom-QAOA ansatz are lower than those for the standard ansatz. For practical purposes, we restrict our cost landscape to the region $\gamma\in[-\pi,\pi]$ and $\beta\in[-\pi/4,\pi/4]$, as optimizing over the entire possible region would be intractable. Also, the global optimum will on average occur near zero for non-periodic functions [19, 4, 5].
@@
One potential concern with our proposed QAOA ansatz is the number of two-qubit gates introduced into the circuit, as two-qubit gates are the primary source of error on modern NISQ devices. The number of two-qubit gates added is equal to the number of new edges in the intersection of $G^{\prime}$ and the complement of $G$. Therefore, the number of additional two-qubit gates depends on the original graph and the method used (e.g., Full or Triangle), providing some flexibility in our approach. Users can adjust the ansatz based on their specific needs. Additionally, if we were to increase $p$ by 1 instead, the number of additional two-qubit gates would equal the number of edges in $G$.
@@
Figure 6: Energy landscapes for (a) standard QAOA and (b) Phantom-QAOA with $G^{\prime}$ constructed using the Triangle method, setting $\alpha=0.7$. The region within the green border is the region we explore in our numerical experiments. We see that greater peak energies can be found outside of the original search space.